### v1

@@ A.1. Datasets
A.2. Implementation Details
1) To keep consistent with previous works, we adopt Mean Squared Error (mse) and Mean Absolute Error (mae) as evaluation metrics. 2) We consider four forecasting horizon $F$: 96, 192, 336, and 720 for the eight common multivariate forecasting datasets. The look-back windows is fixed at 96 for all the baselines. 3) For the 12 real-world datasets that satisfy the TSF-X conditions, we conduct both long-term and short-term prediction experiments. For the Colbun and Raperl datasets, the short-term prediction uses a lookback window of 60 with a prediction horizon of 10, while the long-term prediction uses a lookback window of 180 with a prediction horizon of 30. For the other datasets, the short-term prediction uses a lookback window of 168 with a prediction horizon of 24, and the long-term prediction uses a lookback window of 720 with a prediction horizon of 360. 4) We utilize the TFB (Qiu et al., 2024) code repository for unified evaluation, with all baseline results also derived from TFB. Following the settings in TFB (Qiu et al., 2024) and TSFM-Bench (Li et al., 2025c), we do not apply the “Drop Last” trick to ensure a fair comparison. 5) All experiments of DAG are conducted using PyTorch (Paszke et al., 2019) in Python 3.8 and executed on an NVIDIA Tesla-A800 GPU. The training process is guided by the L1 loss function and employs the ADAM optimizer. The initial batch size is set to 64, with the flexibility to halve it (down to a minimum of 8) in case of an Out-Of-Memory (OOM) issue.
A.3. MLP Fusion Approach
@@
Figure 5. Forecasting performance (MSE) with varying look-back windows on 2 datasets: Electricity and Energy. The look-back windows are selected to be H = 48, 96, 192, 336 and 720, and the forecasting horizons are F = 720 and 360.
@@ Appendix B Experimental Results
B.1. Full Results
Full results on multivariate datasets with TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$) are provided in Table 6.
B.2. Increasing Look-back Window

### v2

@@ A.1 Datasets
A.2 Implementation Details
1) To keep consistent with previous works, we adopt Mean Squared Error (mse) and Mean Absolute Error (mae) as evaluation metrics.
2) For the 12 real-world datasets that satisfy the TSF-X conditions, we conduct both long-term and short-term prediction experiments. For the Colbun and Raperl datasets, the short-term prediction uses a lookback window of 60 with a prediction horizon of 10, while the long-term prediction uses a lookback window of 180 with a prediction horizon of 30. For the other datasets, the short-term prediction uses a lookback window of 168 with a prediction horizon of 24, and the long-term prediction uses a lookback window of 720 with a prediction horizon of 360.
3) We utilize the TFB (Qiu et al., 2024) code repository for unified evaluation, with all baseline results also derived from TFB. Following the settings in TFB (Qiu et al., 2024) and TSFM-Bench (Li et al., 2025), we do not apply the “Drop Last” trick to ensure a fair comparison.
4) All experiments of DAG are conducted using PyTorch (Paszke et al., 2019) in Python 3.8 and executed on an NVIDIA Tesla-A800 GPU. The training process is guided by the L1 loss function and employs the ADAM optimizer. The initial batch size is set to 64, with the flexibility to halve it (down to a minimum of 8) in case of an Out-Of-Memory issue.
A.3 MLP Fusion Approach
@@
Figure 4: Forecasting performance (MSE) with varying look-back windows on 2 datasets: Electricity and Energy. The look-back windows are H = 48, 96, 192, 336 and 720, and the forecasting horizons are F = 720 and 360.
Figure 5: Parameter sensitivity studies of main hyper-parameters in DAG.
@@ Appendix B Experimental Results
B.1 Full Results
Full results on multivariate datasets with TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$) are provided in Table 5. Full results on multivariate datasets, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$) are provided in Table 6.
B.2 Parameter Sensitivity
We also conduct parameter sensitivity studies of DAG. We make the following observations: 1) Figure 5a and Figure 5b illustrate the performance variations of the DAG model under different settings of the fusion weight $\lambda_{1}$ and the correlation weight $\lambda_{2}$. Overall, the model remains stable across a wide range of values, with the optimal range for both $\lambda_{1}$ and $\lambda_{2}$ typically falling between 0.3 and 0.7, indicating that moderate weighting is beneficial for improving performance. 2) Figure 5c shows that model dimension has a certain impact on performance, but DAG maintains good stability across different configurations. Smaller dimensions help reduce computational cost, while dimensions in the range of 64 to 256 generally achieve a good balance between prediction accuracy and model complexity. 3) Figure 5d explores the effect of patch length on model performance. Due to differences in temporal dependencies across datasets, the optimal patch length varies; however, a patch size between 8 and 32 tends to yield favorable results. It is worth noting that very small patch lengths may lead to increased computational complexity, while excessively large patch lengths may weaken the model’s ability to capture local dependencies.
B.3 Increasing Look-back Window
@@
B.4 Visualization of Results
We visualize cases of predictions on the NP dataset in Figure 6, where future exogenous variables are available, and Figure 7, where future exogenous variables are not available. As shown in Figure 6 and Figure 7, DAG consistently outperforms all baselines, both when future exogenous variables are available and when they are not, indicating that it effectively discovers and injects correlation relationships across both temporal and channel dimensions, fully leveraging exogenous variables.
Figure 6: Visualization of prediction results on NP dataset where future exogenous variables are available.
Figure 7: Visualization of prediction results on NP dataset where future exogenous variables are not available.